OpenStudy (anonymous):

ln(x-2)+ln(2x-3)=2lnx Help please.

OpenStudy (whpalmer4):

\[\ln(x-2) + \ln(2x-3) = 2\ln x\] Adding logs is same as taking the log of the product, so: \[\ln((x-2)(2x-3)) = 2\ln x\]Take anti-log of both sides \[(x-2)(2x-3) = x^2\] Now solve that for \(x\). You'll get two solutions, one of which is extraneous because it implies logarithms of negative numbers...

OpenStudy (anonymous):

Thank you!

OpenStudy (whpalmer4):

What did you get for an answer?